# An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function

### v1

@@ 2.5 Manifold learning
Manifold learning aims at learning a norm-preserving and neighborhood-preserving mapping of high-dimensional inputs into a lower-dimensional space [24]. Given a collection of points $\{z_{1},\dots,z_{m}\}$ in $\mathbb{R}^{N}$, Saul [33] looks for a faithful representation $\{y_{1},\dots,y_{m}\}$ in $\mathbb{R}^{r}$, with $r<N$. Let $\tau\in(0,1)$, Saul defines an embedding to be $\tau$-faithful if
@@
Equation (7) implies that the embedding must preserve norms (take $i=j$ to obtain $(1-\tau)\lVert y_{i}\rVert^{2}$ $=$ $(1-\tau)\lVert z_{i}\rVert^{2}$) as well as the small angles between two points $(z_{i},z_{j})$: when $\cos(z_{i},z_{j})>\tau$, we need $\cos(z_{i},z_{j})=\cos(y_{i},y_{j})$. Large angles have to remain large but do not need to be preserved, that is, $\cos(z_{i},z_{j})\leq\tau$ only requires $\cos(y_{i},y_{j})\leq\tau$. The condition (7) defines a similarity between points based on norm and angles, which differs substantially from the more common concept of pairwise distances. This makes it suitable, for example, in text analysis where the $z_{i}$’s are sparse vectors of word counts.
@@
Second, compute a ReLU-decomposition, $X\approx\max(0,\Theta)$, and the matrix $\max(0,\Theta)$ represents the similarities between points in the lower-dimensional space. Finally, the embedded points are recovered from the low-rank approximation $\Theta$; see [33] for more details.
3 ReLU-NMD vs Latent-ReLU-NMD
In this section, we explore the link between ReLU-NMD in (1) and Latent-ReLU-NMD in (2). In particular, we aim to determine if the optimal low-rank approximations provided by the two models coincide. First, when $X$ contains negative values, Latent-ReLU-NMD is infeasible, since there exists no latent variable $Z$ such that $X=\max(0,Z)$. In contrast, ReLU-NMD still admits feasible solutions. Of course, the ReLU decomposition is more meaningful when $X$ is nonnegative which is the case we focus on in the following.
In the exact case, when there exists a low-rank matrix $\Theta$ such that $X=\max(0,\Theta)$, ReLU-NMD and Latent-ReLU-NMD are equivalent: for any low-rank solution $\Theta^{*}$ of (1), there exists a latent matrix $Z^{*}$, such that $(Z^{*},\Theta^{*})$ is a solution of (2) and vice versa. Indeed, in the exact case there always exists a trivial latent solution $Z^{*}=\Theta^{*}$. We show in what follows that this is not true in general, when an exact decomposition does not exist.
We demonstrate first that there exists a specific matrix $X$ such that the rank-1 Latent-ReLU-NMD is ill-posed since the infimum of (2) is not attained (see Lemma 3.1). Moreover, by showing that for the same matrix $X$, ReLU-NMD admits a non-empty set of optimal solutions (Lemma 3.3), we establish that the two formulations do not necessarily yield the same optimal approximations. Note that we could not establish the general well-poseness of ReLU-NMD, which is a direction of future research.
Lemma 3.1.
@@ 3 ReLU-NMD vs Latent-ReLU-NMD
Then the infimum of Latent-ReLU-NMD in (2) is not attained.
@@
We need to show that there exists no solution of Latent-ReLU-NMD attaining the infimum of (2). Let us prove the result by contradiction, and assume there exists a solution to (2), denoted $(\widehat{Z},\widehat{\Theta})$. First note that, given $\widehat{Z}$, the optimal $\widehat{\Theta}$ must be a best rank-1 approximation of $\widehat{Z}$. Hence $\widehat{Z}$ is such that the distance with its best rank-$1$ approximation is minimized and such that it satisfies the constraint $\max(0,\widehat{Z})=X$. Let us denote the $i$-th singular values of $Z$ as $\sigma_{i}(Z)$, and $\widehat{Z}_{1}$ be a best rank-one approximation of $\widehat{Z}$. By the Eckart-Young theorem, $\|\widehat{Z}-\widehat{Z}_{1}\|_{F}^{2}=\sigma_{2}(Z)^{2}$ since $\widehat{Z}$ is a 2-by-2 matrix. Hence $\widehat{Z}$ must belong to
@@
Equivalently, we want to find $b\leq 0$ such that the matrix $Z=\begin{pmatrix}1&b\\ \epsilon&1\end{pmatrix}$ has the lowest possible second singular value. Since $Z$ is a $2\times 2$ matrix, we can compute the eigenvalues of $ZZ^{T}$, solve det$(ZZ^{T}-\sigma I)=0$ and get an explicit expression for the second singular value of $Z$: for a general matrix $\begin{pmatrix}a&b\\ c&d\end{pmatrix},$ the second singular value is given by
$\sigma_{2}=\sqrt{\frac{S_{1}-S_{2}}{2}},$
where
$S_{1}=a^{2}+b^{2}+c^{2}+d^{2},\quad S_{2}=\sqrt{(a^{2}+b^{2}-c^{2}-d^{2})^{2}+4(ac+bd)^{2}}.$
We rewrite (9) as
@@
We demonstrate in the following lemma that rank-1 ReLU-NMD admits an optimal solution for $X$ as in (8).
Lemma 3.3.
Let $X$ as in (8) and $r=1$, then for every $0<\epsilon<\frac{1}{\sqrt{2}}$, the infimum of ReLU-NMD is attained by all the matrices of the form
@@
Let us denote the error function of ReLU-NMD as $F(\Theta)=\lVert X-\max(0,\Theta)\rVert_{F}^{2}$. Note that $2\times 2$ rank-1 matrices have limited patterns that the signs of their entries must follow. For instance, if the matrix does not contain zeros, the number of positive and negative entries must be even, otherwise the rank cannot be equal to one. We proceed by analyzing three different cases.
@@
Case 3: The rank-1 matrix $\Theta$ contains only positive entries. If $\Theta\geq 0$, ReLU-NMD corresponds to the following problem without nonlinear activations:
$\operatorname{argmin}_{\Theta}\lVert X-\Theta\rVert_{F}^{2}\ \text{ such that }\ \Theta\geq 0.$
Therefore, we need to show that there exists no rank-1 matrix with nonnegative entries that achieves a residual value $F(\Theta)$ strictly smaller than $\epsilon^{2}$. Let $\widehat{\sigma}_{2}$ be the second singular value of $X$ as in (8). If rank($\Theta$)=1, then $\lVert X-\Theta\rVert_{F}^{2}\geq\widehat{\sigma}_{2}^{2}$ and using equation (10),
$\widehat{\sigma}_{2}^{2}=\frac{2+\epsilon^{2}-\epsilon\sqrt{\epsilon^{2}+4}}{2}.$
Simple calculations show that $\widehat{\sigma}_{2}^{2}$ is strictly larger than $\epsilon^{2}$ for any $0<\epsilon<\frac{1}{\sqrt{2}}$.
As a consequence of Lemma 3.1 and Lemma 3.3, we have the following corollary.
@@
ReLU-NMD and Latent-ReLU-NMD may not share the same set of optimal low-rank solutions.
We established the non equivalence between the low-rank solutions of ReLU-NMD and Latent-ReLU-NMD. However, in practice, we are interested in computing accurate solutions of ReLU-NMD even though they might not be optimal. Numerous numerical results show that the ReLU decomposition obtained by Latent-ReLU-NMD achieves a small residual error also for ReLU-NMD [34, 35, 41]. The following theorem formalizes this observation, showing that any feasible point $(\widehat{\Theta},\widehat{Z})$ of the Latent model has a residual value for $\widehat{\Theta}$ in the original ReLU-NMD that is upper bounded by twice the residual of Latent-ReLU-NMD.
@@
Let $X\in\mathbb{R}^{n\times m}$ be a nonnegative matrix and denote as $(\widehat{\Theta},\widehat{Z})$ a feasible point of the Latent-ReLU-NMD model in (2), then
(12) $\lVert X-\max(0,\widehat{\Theta})\rVert_{F}^{2}\leq 2\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
@@
We have
(13) $\lVert X-\max(0,\widehat{\Theta})\rVert_{F}^{2}$ $=\lVert X-\widehat{Z}+\widehat{Z}-\max(0,\widehat{\Theta})-\min(0,\widehat{\Theta})+\min(0,\widehat{\Theta})\rVert_{F}^{2}$
$\leq\lVert X-\widehat{Z}+\min(0,\widehat{\Theta})\rVert_{F}^{2}+\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
Let us define $\Omega=\{(i,j)\ |\ X(i,j)>0\}$ and $\widehat{\Omega}=\{(i,j)\ |\ \widehat{\Theta}(i,j)>0\}$, and $\Omega^{C}$ and $\widehat{\Omega}^{C}$ their respective complements. By assumption $\widehat{Z}$ is a feasible solution of (2) and hence $\max(0,\widehat{Z})=X$, therefore
$\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}=$ $\sum_{(i,j)\in\Omega\cap\widehat{\Omega}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+$
$+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}.$
Similarly, observe that $Z_{i,j}\leq 0$ whenever $X_{i,j}=0$, and hence
(14) $\lVert X-\widehat{Z}+\min(0,\widehat{\Theta})\rVert_{F}^{2}=$
$=\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}\widehat{\Theta}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}\widehat{Z}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}$
$\leq\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}$
$\leq\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}^{2}.$
The first inequality holds because
$\widehat{\Theta}_{ij}^{2}\leq(X_{ij}-\widehat{\Theta}_{ij})^{2}\mbox{ for }(i,j)\in\Omega\cap\widehat{\Omega}^{C},\text{ and }$
$\widehat{Z}_{ij}^{2}\leq(\widehat{Z}_{ij}-\widehat{\Theta}_{ij})^{2}\mbox{ for }(i,j)\in\Omega^{C}\cap\widehat{\Omega}.$
@@
A straightforward consequence of Theorem 3.6 is that a solution of Latent-ReLU-NMD with small error provides a ReLU-NMD solution with small error, even though the two models do not necessarily yield the same approximation.
4 Algorithms for the ReLU decomposition
This section is dedicated to the algorithms for computing ReLU decompositions and to the study of their convergence. In Section 4.1, we briefly review the state-of-the-art algorithms for solving Latent-ReLU-NMD and 3B-ReLU-NMD. In Section 4.2, we discuss the convergence of the BCD-NMD scheme. Finally, we present our new algorithm eBCD-NMD along with its convergence in Section 4.3.

### v2

@@ 2.4 Manifold learning
Manifold learning aims at learning a norm-preserving and neighborhood-preserving mapping of high-dimensional inputs into a lower-dimensional space [lee2007nonlinear]. Given a collection of points $\{z_{1},\dots,z_{m}\}$ in $\mathbb{R}^{N}$, Saul [saul2022geometrical] looks for a faithful representation $\{y_{1},\dots,y_{m}\}$ in $\mathbb{R}^{r}$, with $r<N$. Let $\tau\in(0,1)$, an embedding is $\tau$-faithful if
@@
Equation (6) implies that the embedding must preserve norms (take $i=j$ to obtain $(1-\tau)\lVert y_{i}\rVert^{2}$ == $(1-\tau)\lVert z_{i}\rVert^{2}$) as well as the small angles between two points $(z_{i},z_{j})$: when $\cos(z_{i},z_{j})>\tau$, we need $\cos(z_{i},z_{j})=\cos(y_{i},y_{j})$. Large angles have to remain large but do not need to be preserved, that is, $\cos(z_{i},z_{j})\leq\tau$ only requires $\cos(y_{i},y_{j})\leq\tau$. The condition (6) defines a similarity between points based on norm and angles, which differs substantially from the more common concept of pairwise distances. This makes it suitable, for example, in text analysis where the $z_{i}$’s are sparse vectors of word counts.
@@
Second, compute an RMD, $X\approx\max(0,\Theta)$, and the matrix $\max(0,\Theta)$ represents the similarities between points in the lower-dimensional space. Finally, the embedded points are recovered from the low-rank approximation $\Theta$; see [saul2022geometrical] for more details.
3 LS-RMD vs Latent-RMD
In this section, we explore the link between LS-RMD in (1) and Latent-RMD in (2). In particular, we aim to determine if the optimal low-rank approximations provided by the two models coincide. First, when $X$ contains negative values, Latent-RMD is infeasible since there exists no latent variable $Z$ such that $X=\max(0,Z)$. In contrast, LS-RMD still admits feasible solutions. Of course, the RMD is more meaningful when $X$ is nonnegative, which is the case we focus on in the following.
In the exact case, if there exists a low-rank matrix $\Theta$ such that $X=\max(0,\Theta)$, LS-RMD and Latent-RMD are equivalent: for any low-rank solution $\Theta^{*}$ of (1), there exists a latent matrix $Z^{*}$, such that $(Z^{*},\Theta^{*})$ is a solution of (2) and vice versa. Indeed, in this case, there always exists a trivial latent solution $Z^{*}=\Theta^{*}$. We show in what follows that this is not always true if an exact decomposition does not exist.
We demonstrate first that there exists a specific matrix $X$ such that the rank-1 Latent-RMD is ill-posed since the infimum of (2) is not attained (see Example 3.1). Moreover, by showing that for the same matrix $X$, LS-RMD admits a non-empty set of optimal solutions ( Example 3.3), we establish that the two formulations do not necessarily yield the same optimal approximations. Note that we could not establish the general well-poseness of LS-RMD, which is a direction of future research.
Example 3.1.
@@ 3 LS-RMD vs Latent-RMD
Then the infimum of Latent-RMD in (2) is not attained.
@@
[…]
@@
[…]
@@
We now demonstrate that rank-1 LS-RMD admits an optimal solution for $X$ as in (7).
Example 3.3.
Let $X$ as in (7) and $r=1$, then for every $0<\epsilon<\frac{1}{\sqrt{2}}$, the infimum of LS-RMD is attained by all the matrices of the form
@@
Let us denote the error function of LS-RMD as $F(\Theta)=\lVert X-\max(0,\Theta)\rVert_{F}^{2}$. Note that $2\times 2$ rank-1 matrices have limited patterns that the signs of their entries must follow. For instance, if the matrix does not contain zeros, the number of positive and negative entries must be even; otherwise, the rank cannot be equal to one. We proceed by analyzing three different cases.
@@
Case 3: The rank-1 matrix $\Theta$ contains only positive entries. If $\Theta\geq 0$, the nonlinear activation has no effect on $\Theta$ and the ReLU-NMD corresponds to minimizing the function $\lVert X-\Theta\rVert_{F}^{2}$. Let $\widehat{\sigma}_{2}$ be the second singular value of $X$ as in (7). If rank($\Theta$)=1, then $\lVert X-\Theta\rVert_{F}^{2}\geq\widehat{\sigma}_{2}^{2}$ and using equation (9), we have $\widehat{\sigma}_{2}^{2}=\frac{2+\epsilon^{2}-\epsilon\sqrt{\epsilon^{2}+4}}{2}.$ Simple calculations show that $\widehat{\sigma}_{2}^{2}$ is strictly larger than $\epsilon^{2}$ for any $0<\epsilon<\frac{1}{\sqrt{2}}$.
As a consequence of Example 3.1 and Example 3.3, we have the following corollary.
@@
LS-RMD and Latent-RMD may not share the same set of optimal low-rank solutions.
We established the non equivalence between the low-rank solutions of LS-RMD and Latent-RMD. However, in practice, we are interested in computing accurate solutions of even though they might not be optimal. Numerous numerical results show that the decompositions obtained by solving the Latent-RMD model achieves a small residual error also for LS-RMD [saul2022nonlinear, seraghiti2023accelerated, wang2024momentum]. The following theorem formalizes this observation, showing that any feasible point $(\bar{\Theta},\bar{Z})$ of the latent model has a residual value for in the original LS-RMD that is upper bounded by four times the residual of Latent-RMD.
@@
Let $X\in\mathbb{R}^{n\times m}$ be a nonnegative matrix and denote as $(\bar{\Theta},\bar{Z})$ a feasible point of the Latent-RMD model in (2); then
(11) $\lVert X-\max(0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}})\rVert_{F}^{2}\leq{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}}\lVert{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}\rVert_{F}^{2}$.
@@
Using the matrix inequality $\lVert A+B\rVert_{F}^{2}\leq 2\lVert A\rVert_{F}^{2}+2\lVert B\rVert_{F}^{2}$, we have
(12) $\lVert X-\max(0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}})\rVert_{F}^{2}$ $=\lVert X-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\min(0,\bar{\Theta})-\min(0,\bar{\Theta})-\max(0,\bar{\Theta})}}\rVert_{F}^{2}$
[…]
$+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}.$
Similarly, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{i,j}\leq 0$ whenever $X_{i,j}=0$; thus,
$\lVert X-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}+\min(0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}})\rVert_{F}^{2}$ $=\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}$
$\leq\sum_{(i,j)\in\Omega\cap\widehat{\Omega}^{C}}(X_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}+\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}+\hskip-5.69046pt\sum_{(i,j)\in\Omega^{C}\cap\widehat{\Omega}^{C}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}$
[…]
where the first inequality holds because
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij}^{2}\leq(X_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}\mbox{ for }(i,j)\in\Omega\cap\widehat{\Omega}^{C},\quad\text{ and }\quad{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}^{2}\leq({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{Z}}}_{ij}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Theta}}}_{ij})^{2}\mbox{ for }(i,j)\in\Omega^{C}\cap\widehat{\Omega}$.
@@
A straightforward consequence of Theorem 3.6 is that a solution of Latent-RMD with a small error provides a LS-RMD solution with a small error, even though the two models do not necessarily yield the same approximation. We specify that the same results holds if we reparametrize $\Theta$ as $WH$; thus, the same bound holds for 3B-RMD as well.
4 Algorithms for RMD
This section is dedicated to the algorithms for computing RMDs and to the study of their convergence. In Section 4.1, we briefly review the state-of-the-art algorithms for solving Latent-RMD and 3B-RMD. In Section 4.2, we discuss the convergence of the BCD scheme. Finally, we present our new algorithm eBCD along with its convergence in Section 4.3.